# DenseSplat: Densifying Gaussian Splatting SLAM with Neural Radiance Prior

### v1

@@ III-C Loop Closure and Bundle Adjustment
In the bundle adjustment (BA) process, we use the Bag of Words (BoW) model [31] to determine the relevance between keyframes. Upon detecting a loop, this triggers a BA procedure for the involved keyframes, akin to the methods described in [50, 53].
To maintain the geometric and visual consistency of the map after the BA process, we adjust the rendered color $\mathcal{R}(\mathbf{T}_{j},\mathcal{G}_{\mathbf{c}})$ and depth $\mathcal{R}(\mathbf{T}_{j},\mathcal{G}_{\mathbf{d}})$ to warp the co-visible keyframe $k$ using the estimated relative pose transformation $\mathbf{T}_{j}^{k}$. We then construct the BA loss using the following equations:
$\mathcal{L}_{\rm BA}^{\mathbf{C}}=\sum_{k=1}^{N-1}\sum_{j=k+1}^{N}\|\mathbf{T}_{j}^{k}\cdot\mathcal{R}(\mathbf{T}_{j},\mathcal{G}_{\mathbf{c}})-\mathcal{R}(\mathbf{T}_{k},\mathcal{G}_{\mathbf{c}})\|$ (9)
$\mathcal{L}_{\rm BA}^{\mathbf{D}}=\sum_{k=1}^{N-1}\sum_{j=k+1}^{N}\|\mathbf{T}_{j}^{k}\cdot\mathcal{R}(\mathbf{T}_{j},\mathcal{G}_{\mathbf{d}})-\mathcal{R}(\mathbf{T}_{k},\mathcal{G}_{\mathbf{d}})$ (10)
Leveraging the rapid rendering capabilities of 3D Gaussian Splatting, our method supports real-time re-rendering and the correction of BA-induced drift errors.
@@ III-D Gaussian Map Optimization
where $\mathbf{C}_{\rm GT}$ and $\mathbf{D}_{\rm GT}$ denote the ground-truth color and depth from the input stream. The SSIM loss [64] calculates the structural similarity between the rendered and ground-truth images. The coefficients $\lambda_{\rm c}$, $\lambda_{\rm d}$, $\lambda_{\rm ssim}$, and $\lambda_{\rm reg}$ are weighting hyperparameters.
TABLE I: Quantitative comparison of our method and the baselines in training view rendering on the Replica dataset [65]. The Underline indicates that relocalization was triggered due to accumulated tracking errors in the ORB tracking system [31]. Dash presents system failure.
@@
Submap Division and Fusion When deploying a SLAM system in large-scale environments, managing the excessive memory consumption of the dense mapping is critical for practical application. To address this issue, we employ a simple yet effective submap division and fusion strategy as presented in Algorithm 1. Specifically, we partition the input frames into submaps at intervals of every 400 frames, structured as follows:
$\left\{I_{i},D_{i}\right\}_{i=1}^{N}\mapsto\left\{\mathrm{SF}_{\langle f_{\theta},\mathcal{G}\rangle}^{1},\mathrm{SF}_{\langle f_{\theta},\mathcal{G}\rangle}^{2},\ldots,\mathrm{SF}_{\langle f_{\theta},\mathcal{G}\rangle}^{n}\right\}~{},$ (12)
where $\mathrm{SF}_{\langle f_{\theta},\mathcal{G}\rangle}^{n}$ represents each submap used to develop NeRF models and subsequent Gaussian maps. While the explicit Gaussian representation enables the seamless combination of submaps into a global map, directly fusing submaps remains a challenging task. Drawing inspiration from Mipsfusion [66], we utilize anchor-frame BA during our submap fusion process to achieve precise alignment and seamless fusion at submap boundaries. Each submap is anchored based on the estimated pose of its first frame. Following BA, we precisely adjust the central pose of each submap to ensure accurate re-anchoring. Our submap strategy effectively reduces memory consumption by enabling the parallel construction of each submap, thereby mitigating the issues associated with the continuous expansion of the global map.
@@
9:     if CheckKeyframe($\{I_{i},D_{i},\mathbf{T}_{i}\}$then
@@
18:     if Length($\Lambda$) $>j$ then
@@ IV Experiments
Figure 3: Novel-view synthesis (NVS) comparison of DenseSplat and Gaussian-based baseline methods on the selected scenes of Replica apartment dataset [65]. Our method demonstrates superior performance in geometric accuracy, hole filling, and fine-grained texture rendering. Crucially, DenseSplat utilizes a sparse keyframe (kf) interval of 20, offering a more efficient and practical setup compared to the dense keyframe lists employed by the baseline methods.
@@
Implementation Details The experiments were conducted on a single NVIDIA A100 GPU with 80 GB of VRAM. The keyframe interval was set to 20. The tracking and mapping iterations for the NeRF model are both set to 10, and the mapping iteration for the Gaussian map is 30. Parameters of the NeRF model and Gaussian primitives were optimized using the Adam optimizer, with the learning rate following the settings in [60] and [3]. The initial pose is estimated purely using NeRF with a tracking iteration of 10. For mapping, NeRF undergoes 10 iterations, while the Gaussian map is optimized over 30 iterations for fine-grained reconstruction. The keyframe intervals for both NeRF and Gaussian models are 20 frames for the Replica [65] and TUM RGB-D [67] datasets, and 25 frames for the ScanNet dataset [68]. For NeRF, we employ a sampling ratio of 64 in our geometry-warping initialization. The density threshold $\tau_{\rm grid}$ for selecting transmission points is set at 0.001. In the ray-guided pruning process, the pruning threshold $\tau_{\rm prune}$ is set at 0.001 to optimize memory and computational resources. During map optimization, the regularization scale threshold $\tau_{\rm scale}$ is maintained at 1.0, where scales exceeding this threshold are penalized. The weighting hyperparameters $\lambda_{\rm c}$, $\lambda_{\rm d}$, $\lambda_{\rm ssim}$, and $\lambda_{\rm reg}$ are set at 0.8, 0.1, 0.2, and 0.001 respectively, promoting consistency across various metrics including color accuracy, depth precision, structural similarity, and regularization.
@@
Datasets and Metrics We conduct comprehensive evaluations using both synthesized and real-world scenes from Replica [65], ScanNet [68], and TUM RGB-D [67] datasets (supplementary material). In addition, we use four large-scale apartment scenes from the Replica dataset [65], with the largest scene containing up to two floors and eight rooms, presenting challenging indoor layouts with complex corridors and stairs. To evaluate tracking accuracy, we employ ATE RMSE (cm). For reconstruction results on training views, we adhere to PSNR, SSIM, and LPIPS to quantitatively evaluate rendering quality. To qualitatively compare mapping quality, we visualize the reconstruction outcomes from novel views. The running speed and computation usage are assessed using FPS and GPU consumption. Best results are shaded as first, second, and third.
@@
Baseline Methods We compare DenseSplat with the NeRF-based dense RGB-D SLAM systems, including Co-SLAM [48], Point-SLAM [49], Loopy-SLAM [50]; as well as recent 3DGS-based systems such as SplaTAM [18], MonoGS [19], Gaussian-SLAM [52], Photo-SLAM [20], and RTG-SLAM [21]. To ensure a fair comparison, all results are assessed based on the final global maps produced by the SLAM systems.
@@ IV-A Evaluation of Tracking and Mapping
TABLE II: Evaluation of averaged tracking performance, system frame rate, memory consumption, and map size across the eight scenes of the Replica dataset [65]. We separately present the systems that utilize ORB-SLAM [29, 31] and frame-to-frame tracking.
[…]
ORB Photo-SLAM [20] 0.59 41.64 30.36 20.71 8.02 59
RTG-SLAM [21] 0.49 50.33 20.09 17.30 10.18 71
Co-SLAM [48] 1.12 10.2 10.0 9.26 7.90 7
Frame-to-frame Point-SLAM [49] 0.54 0.95 0.87 0.44 18.86 154
Loopy-SLAM [50] 0.29 0.95 0.87 0.43 18.91 177
SplaTAM [18] 0.55 0.86 0.51 0.51 11.27 331
MonoGS [19] 0.58 3.41 3.16 3.09 13.88 42
Ours 0.33 7.96 23.35 7.67 6.67 117
@@
Evaluations on Replica Dataset We quantitatively assessed the rendering quality of our method against NeRF-based and Gaussian-based SLAM systems on the Replica dataset [65], as presented in Table I. DenseSplat exhibits competitive rendering quality in single-room scale scenes, achieving state-of-the-art performance in some instances. Moreover, our method substantially outperforms baseline approaches in large-scale, multi-room environments, such as apartment scenes. Specifically, Gaussian SLAM baselines struggle due to ineffective bundle adjustment and cumbersome map representations, which do not adequately address the extensive camera movement and complex geometry characteristic of large-scale scenes. NeRF baselines provide robust tracking and efficient implicit scene representation but do not match our method in detailed reconstruction and rendering quality. DenseSplat integrates the strengths of both NeRF and Gaussian maps, utilizing the robust NeRF prior to stabilize scene representation. In addition, our submap strategy facilitates parallel computation of maps, significantly reducing memory consumption that often leads to system failures in Gaussian systems, such as SplaTAM [18] in apartment_0. The novel-view rendering results, shown in Figure 3, qualitatively compare our reconstructed map with Gaussian-based systems. DenseSplat demonstrates superior mapping quality compared to baselines that suffer from significant scene drift and floaters.
@@
To assess tracking accuracy and system efficiency, we compared the average ATE and frame rates of each component as detailed in Table II. DenseSplat delivers competitive tracking performance while achieving the highest mapping frame rate among frame-to-frame systems. Additionally, our map division strategy results in the lowest memory consumption during runtime compared to the baseline methods.
@@
TABLE III: Quantitative comparison of our method and the baselines in training view rendering on the ScanNet dataset [68]. Our method demonstrates SOTA performances among three metrics.
  Methods Metrics $0$ 0059 0106 0169 0181 0207 Avg.
@@
Evaluations on ScanNet Dataset We present quantitative evaluations of our method alongside the baseline approaches using 6 scenes from the ScanNet dataset [68]. As shown in Table III, we evaluate rendering quality by comparing PSNR, SSIM, and LPIPS metrics across training views. In real-world environments, challenges on noisy input such as depth map errors and blurred images introduce significant difficulties for map initialization and optimization. Additionally, recent Gaussian-based methods like SplaTAM [18] and MonoGS [19], which do not incorporate loop-closure, experience considerable loop-induced map drift that remarkably reduces reconstruction quality. In contrast, our method utilizes the robust NeRF prior and integrates both the bundle adjustment (BA) and subsequent BA-induced map refinement modules, achieving state-of-the-art performance. Table IV also presents the tracking performance using ATE RMSE (cm) on the ScanNet dataset [68]. Our method demonstrates superior tracking accuracy compared to Gaussian baseline methods [18, 19] that do not incorporate the BA process. When comparing with Photo-SLAM [20] and RTG-SLAM [21], which incorporate point-based tracking systems derived from ORB-SLAM [29, 31], DenseSplat achieves lower ATE errors, benefiting from its robust tracking based on the NeRF model.
TABLE IV: The performance of ATE RMSE (cm) on 6 scenes from the ScanNet [68] dataset. For our method, we provide results both with and without Bundle Adjustment (BA).
Methods $0$ 0059 0106 0169 0181 0207 Avg.
ORB Photo-SLAM [20] 7.62 7.94 9.36 10.01 22.97 7.23 10.85
RTG-SLAM [21] 8.04 6.82 9.22 10.15 24.36 9.25 11.31
ESLAM [47] 7.54 8.52 7.39 8.17 9.13 5.61 7.73
Co-SLAM [48] 7.13 11.14 9.36 5.90 11.81 7.14 8.75
Frame-to-frame Loopy-SLAM [50] 4.28 7.59 8.37 7.56 10.68 7.95 7.70
SplaTAM [18] 12.83 10.10 17.72 12.08 11.10 7.47 11.88
MonoGS [19] 15.94 6.41 19.44 10.44 12.23 10.46 12.48
@@
Figure 4 visualizes the novel-view synthesis of representative scenes from the ScanNet dataset [68]. DenseSplat offers superior geometric accuracy, for instance, accurately capturing the details of the bicycle in scene_0000, which often suffers from scene drift due to trajectory loops. Compared to Photo-SLAM [20], our method demonstrates comparable tracking performance while providing more fine-grained and complete mapping.
TABLE V: The performance of ATE RMSE (cm) on the TUM RGB-D dataset [67]. For our method, we provide results both with and without Bundle Adjustment (BA).
@@
ORB Photo-SLAM [20] 2.61 0.35 1.00 1.32
RTG-SLAM [21] 1.66 0.38 1.13 1.06
ESLAM [47] 2.30 1.10 2.40 2.00
Co-SLAM [48] 2.70 1.90 2.60 2.40
Frame-to-frame Loopy-SLAM [50] 3.79 1.62 3.41 2.94
SplaTAM [18] 3.35 1.24 5.16 3.25
MonoGS [19] 1.52 1.58 1.65 1.58
@@
Figure 4: NVS comparison on real-world ScanNet dataset [68]. DenseSplat shows superior geometry accuracy and hole filling.
Evaluations on TUM RGB-D Dataset We also conducted experiments on the TUM RGB-D dataset [67] and compared the ATE RMSE (cm) with both feature-based and frame-to-frame tracking systems in Table V. Our system achieves performance comparable to current state-of-the-art methods.
Figure 5: Visualization of NVS of the Replica dataset [65]. The rendered views are shown along with plots for the center of primitives, depicted as blue dots of the same size across all methods. For our method, we explicitly present the NVS at keyframe intervals (kf) at 4 and 20. Looking at objects like floors and tables, DenseSplat demonstrates more evenly distributed primitives compared to baseline methods.

### v2

@@ III-C Loop Closure and Bundle Adjustment
In the bundle adjustment (BA) process, we use the Bag of Words (BoW) model [campos2021orb] to determine the relevance between keyframes. Upon detecting a loop, this triggers a BA procedure for the involved keyframes, in a manner akin to [liso2024loopy, zhu2024loopsplat]. The implementation is detailed in Algorithm 1.
Algorithm 1 Loop Closure and Bundle Adjustment
1:Input: Keyframe stream $\{I_{k},D_{k},\mathbf{T}_{k}\}$, BoW database $\mathcal{B}$
2:Output: Updated keyframe poses $\{\mathbf{T}_{k}\}$
3:for each new keyframe $(I_{k},D_{k},\mathbf{T}_{k})$ do
4:  // Add keyframe to BoW database
5:  $\mathcal{B}\leftarrow\mathcal{B}\cup\{I_{k}\}$
6:  // Retrieve loop closure candidates
7:  $\{j\}\leftarrow\text{QueryBoW}(\mathcal{B},I_{k})$
8:  for each candidate $j$ do
9:   if LoopGeometricCheck($I_{k},I_{j}$) then
10:     // Loop closure confirmed
11:     $\mathbf{T}_{k\rightarrow j}\leftarrow\text{EstimateRelativePose}(I_{k},I_{j})$
12:     // Collect covisible keyframes
13:     $\mathcal{K}_{\text{local}}\leftarrow\text{GetCovisibleKeyframes}(k,j)$
14:     // Pose-only Bundle Adjustment
15:     $\{\mathbf{T}_{p}\}\leftarrow\text{RunPoseBA}(\mathcal{K}_{\text{local}})$
16:     // Update keyframe poses
17:     for each $p\in\mathcal{K}_{\text{local}}$ do
18:      $\mathbf{T}_{p}\leftarrow\text{OptimizedPose}(p)$
19:     end for
20:   end if
21:  end for
22:end for
While BA refines the camera trajectories, the dense map cannot trivially incorporate these updates and remains synchronized with the pre-optimized poses, which leads to spatial drift. To maintain the geometric and visual consistency of the map after the BA process, we adjust the color $\mathbf{C}_{\mathcal{G}}\langle\mathbf{T}_{i}\rangle$ and depth $\mathbf{D}_{\mathcal{G}}\langle\mathbf{T}_{i}\rangle$ rendered at current pose $\mathbf{T}_{i}$ to warp the co-visible keyframe pose $\mathbf{T}_{k}$ using the estimated relative pose transformation $\langle\mathbf{T}_{i\rightarrow k}\rangle$. We then construct the BA-induced mapping loss using the following equations:
$\mathcal{L}_{\rm BA}^{\mathbf{C}}$ $=\sum_{k=1}^{N-1}\sum_{i=k+1}^{N}\|\mathbf{C}_{\mathcal{G}}\langle\mathbf{T}_{i\rightarrow k}\rangle-\mathbf{C}_{\mathcal{G}}\langle\mathbf{T}_{k}\rangle\|$ (9)
$\mathcal{L}_{\rm BA}^{\mathbf{D}}$ $=\sum_{k=1}^{N-1}\sum_{i=k+1}^{N}\|\mathbf{D}_{\mathcal{G}}\langle\mathbf{T}_{i\rightarrow k}\rangle-\mathbf{D}_{\mathcal{G}}\langle\mathbf{T}_{k}\rangle\|$ (10)
Leveraging the rapid rendering capabilities of 3D Gaussian Splatting, our method supports real-time re-rendering and the correction of drift errors.
@@ III-D Gaussian Map Optimization
where $\mathbf{C}_{\rm GT}$ and $\mathbf{D}_{\rm GT}$ denote the ground-truth color and depth from the input stream. The SSIM loss [wang2004image] calculates the structural similarity between the rendered and ground-truth images. The coefficients $\lambda_{\rm c}$, $\lambda_{\rm d}$, $\lambda_{\rm ssim}$, and $\lambda_{\rm reg}$ are weighting hyperparameters.
TABLE I: Quantitative comparison of our method and the baselines in training view rendering on the Replica dataset [straub2019replica]. The Underline indicates that relocalization was triggered due to accumulated tracking errors in the ORB tracking system [campos2021orb]. Dash presents system failure.
@@
Submap Division and Fusion When deploying a SLAM system in large-scale environments, managing the excessive memory consumption of the dense mapping is critical for practical applications. To address this issue, we employ an effective submap division and fusion strategy presented in Algorithm 2. Specifically, we partition input frames into submaps at intervals of every 400 frames, structured as follows:
$\left\{I_{i},D_{i}\right\}_{i=1}^{N}\mapsto\left\{\mathrm{SF}_{\langle f_{\theta}^{1},\mathcal{G}^{1}\rangle}^{1},\mathrm{SF}_{\langle f_{\theta}^{2},\mathcal{G}^{2}\rangle}^{2},\ldots,\mathrm{SF}_{\langle f_{\theta}^{n},\mathcal{G}^{n}\rangle}^{n}\right\}~,$ (12)
where $\mathrm{SF}_{\langle f_{\theta},\mathcal{G}\rangle}^{n}$ represents each submap used to develop NeRF models and subsequent Gaussian maps. While the explicit Gaussian representation enables the seamless combination of submaps into a global map, directly fusing submaps remains a challenging task. Drawing inspiration from Mipsfusion [tang2023mips], we utilize anchor-frame BA during our submap fusion process to achieve precise alignment and seamless fusion at submap boundaries. Each submap is anchored based on the estimated pose of its first frame. Following BA, we precisely adjust the central pose of each submap to ensure accurate re-anchoring. Our submap strategy effectively reduces memory consumption by enabling the parallel construction of each submap, thereby mitigating the issues associated with the continuous expansion of the global map.
@@
9:  if CheckKeyframe($\{I_{i},D_{i},\mathbf{T}_{i}\}$) then
@@
18:  if Length($\Lambda$) >j>j then
@@ IV Experiments
Figure 4: Novel-view synthesis (NVS) comparison of DenseSplat and Gaussian-based baseline methods on the selected scenes of Replica apartment dataset [straub2019replica]. Our method demonstrates superior performance in geometric accuracy, hole filling, and fine-grained texture rendering. Crucially, DenseSplat utilizes a sparse keyframe (kf) interval of 20, offering a more efficient and practical setup compared to the dense keyframe lists employed by the baseline methods.
@@
Implementation Details The experiments were conducted on a single NVIDIA A100 GPU with 80 GB of VRAM. The keyframe interval was set to 20. The tracking and mapping iterations for the NeRF model are both set to 10, and the mapping iteration for the Gaussian map is 30. Parameters of the NeRF model and Gaussian primitives were optimized using the Adam optimizer, with the learning rate following the settings in [muller2022instant] and [kerbl20233d]. The initial pose is estimated purely using NeRF with a tracking iteration of 10. For mapping, NeRF undergoes 10 iterations, while the Gaussian map is optimized over 30 iterations for fine-grained reconstruction. The keyframe intervals for both NeRF and Gaussian models are 20 frames for the Replica [straub2019replica] and TUM RGB-D [sturm2012evaluating] datasets, and 25 frames for the ScanNet dataset [dai2017scannet]. For NeRF, we employ a sampling ratio of 64 in our geometry-warping initialization. The density threshold $\tau_{\rm grid}$ for selecting transmission points is set at 0.001. In the ray-guided pruning process, the pruning threshold $\tau_{\rm prune}$ is set at 0.001 to optimize memory and computational resources. During map optimization, the regularization scale threshold $\tau_{\rm scale}$ is maintained at 1.0, where scales exceeding this threshold are penalized. The weighting hyperparameters $\lambda_{\rm c}$, $\lambda_{\rm d}$, $\lambda_{\rm ssim}$, and $\lambda_{\rm reg}$ are set at 0.8, 0.1, 0.2, and 0.001 respectively, promoting consistency across various metrics including color accuracy, depth precision, structural similarity, and regularization.
@@
Datasets and Metrics We conduct comprehensive evaluations using both synthesized and real-world scenes from Replica [straub2019replica], ScanNet [dai2017scannet], and TUM RGB-D [sturm2012evaluating] datasets (supplementary material). In addition, we use four large-scale apartment scenes from the Replica dataset [straub2019replica], with the largest scene containing up to two floors and eight rooms, presenting challenging indoor layouts with complex corridors and stairs. To evaluate tracking accuracy, we employ ATE RMSE (cm). For reconstruction results on training views, we adhere to PSNR, SSIM, and LPIPS to quantitatively evaluate rendering quality. To qualitatively compare mapping quality, we visualize the reconstruction outcomes from novel views. The running speed and computation usage are assessed using FPS and GPU consumption. Best results are shaded as first, second, and third.
@@
Baseline Methods We compare DenseSplat with NeRF-based RGB-D SLAM systems, including Co-SLAM [wang2023co], Point-SLAM [sandstrom2023point], Loopy-SLAM [liso2024loopy]; as well as recent 3DGS-based systems such as SplaTAM [keetha2024splatam], MonoGS [matsuki2024gaussian], Gaussian-SLAM [yugay2023gaussian], Photo-SLAM [huang2024photo], and RTG-SLAM [peng2024rtg]. To ensure a fair comparison, all results are assessed based on the final global maps produced by the SLAM systems.
TABLE II: Evaluation of averaged tracking performance, system frame rate, memory consumption, and map size across the eight scenes of the Replica dataset [straub2019replica]. We separately present the systems that utilize ORB-SLAM [mur2017orb, campos2021orb] and frame-to-frame tracking.
[…]
ORB Photo-SLAM [huang2024photo] 0.59 41.64 30.36 20.71 8.02 59
RTG-SLAM [peng2024rtg] 0.49 50.33 20.09 17.30 10.18 71
Co-SLAM [wang2023co] 1.12 10.2 10.0 9.26 7.90 7
Frame-to-frame Point-SLAM [sandstrom2023point] 0.54 0.95 0.87 0.44 18.86 154
Loopy-SLAM [liso2024loopy] 0.29 0.95 0.87 0.43 18.91 177
SplaTAM [keetha2024splatam] 0.55 0.86 0.51 0.51 11.27 331
MonoGS [matsuki2024gaussian] 0.58 3.41 3.16 3.09 13.88 42
Ours 0.33 7.96 23.35 7.67 6.67 117
@@ IV-A Evaluation of Tracking and Mapping
Evaluations on Replica Dataset We quantitatively assessed the rendering quality of our method against NeRF-based and Gaussian-based SLAM systems on the Replica dataset [straub2019replica], as presented in Table I. DenseSplat exhibits competitive rendering quality in single-room scale scenes, achieving state-of-the-art performance in some instances. Moreover, our method substantially outperforms baseline approaches in large-scale, multi-room environments, such as apartment scenes. Specifically, Gaussian SLAM baselines struggle due to ineffective bundle adjustment and cumbersome map representations, which do not adequately address the extensive camera movement and complex geometry characteristic of large-scale scenes. NeRF baselines provide robust tracking and efficient implicit scene representation but do not match our method in detailed reconstruction and rendering quality. DenseSplat integrates the strengths of both NeRF and Gaussian maps, utilizing the robust NeRF prior to stabilize scene representation. In addition, our submap strategy facilitates parallel computation of maps, significantly reducing memory consumption that often leads to system failures in Gaussian systems, such as SplaTAM [keetha2024splatam] in apartment_0. The novel-view rendering results, shown in Figure 4, qualitatively compare our reconstructed map with Gaussian-based systems. DenseSplat demonstrates superior mapping quality compared to baselines that suffer from significant scene drift and floaters.
@@
To assess tracking accuracy and system efficiency, we compared the average ATE and frame rates as detailed in Table II. DenseSplat delivers competitive tracking performance while achieving the highest mapping frame rate among frame-to-frame systems. Additionally, our map division strategy results in the lowest memory consumption during runtime compared to the baseline methods.
Figure 5: NVS comparison on real-world ScanNet dataset [dai2017scannet]. DenseSplat shows superior geometry accuracy and hole filling.
Figure 6: Visualization of NVS of the Replica dataset [straub2019replica]. The rendered views are shown along with plots for the center of primitives, depicted as blue dots of the same size across all methods. For our method, we explicitly present the NVS at keyframe intervals (kf) at 4 and 20. Looking at objects like floors and tables, DenseSplat demonstrates more evenly distributed primitives compared to baseline methods.
Evaluations on ScanNet Dataset We present quantitative evaluations of our method alongside the baseline approaches using 6 scenes from the ScanNet dataset [dai2017scannet]. As shown in Table III, we evaluate rendering quality by comparing PSNR, SSIM, and LPIPS metrics across training views. In real-world environments, challenges on noisy input such as depth map errors and blurred images introduce significant difficulties for map initialization and optimization. Additionally, recent Gaussian-based methods like SplaTAM [keetha2024splatam] and MonoGS [matsuki2024gaussian], which do not incorporate loop-closure, experience considerable loop-induced map drift that remarkably reduces reconstruction quality. In contrast, our method utilizes the robust NeRF prior and integrates both the bundle adjustment (BA) and subsequent BA-induced map refinement modules, achieving state-of-the-art performance. Table IV also presents the tracking performance using ATE RMSE (cm) on the ScanNet dataset [dai2017scannet]. Our method demonstrates superior tracking accuracy compared to Gaussian baseline methods [keetha2024splatam, matsuki2024gaussian] that do not incorporate the BA process. When comparing with Photo-SLAM [huang2024photo] and RTG-SLAM [peng2024rtg], which incorporate point-based tracking systems derived from ORB-SLAM [mur2017orb, campos2021orb], DenseSplat achieves lower ATE errors, benefiting from its robust tracking based on the NeRF model.
@@
TABLE III: Quantitative comparison of our method and the baselines in training view rendering on the ScanNet dataset [dai2017scannet]. Our method demonstrates SOTA performances.
  Methods Metrics 0000 0059 0106 0169 0181 0207 Avg.
@@
TABLE IV: The performance of ATE RMSE (cm) on 6 scenes from the ScanNet [dai2017scannet] dataset. For our method, we provide results both with and without Bundle Adjustment (BA).
Methods 0000 0059 0106 0169 0181 0207 Avg.
ORB Photo-SLAM [huang2024photo] 7.62 7.94 9.36 10.01 22.97 7.23 10.85
RTG-SLAM [peng2024rtg] 8.04 6.82 9.22 10.15 24.36 9.25 11.31
ESLAM [johari2023eslam] 7.54 8.52 7.39 8.17 9.13 5.61 7.73
Co-SLAM [wang2023co] 7.13 11.14 9.36 5.90 11.81 7.14 8.75
Frame-to-frame Loopy-SLAM [liso2024loopy] 4.28 7.59 8.37 7.56 10.68 7.95 7.70
SplaTAM [keetha2024splatam] 12.83 10.10 17.72 12.08 11.10 7.47 11.88
MonoGS [matsuki2024gaussian] 15.94 6.41 19.44 10.44 12.23 10.46 12.48
@@
Figure 5 shows the novel-view synthesis of representative scenes from the ScanNet dataset [dai2017scannet]. DenseSplat offers superior geometric accuracy, for instance, accurately capturing the details of the bicycle in scene_0000, which often suffers from scene drift due to trajectory loops. Compared to Photo-SLAM [huang2024photo], our method demonstrates comparable tracking performance while providing more fine-grained and complete map reconstruction.
Figure 7: Visualization of hole-filling in novel-view synthesis for the scene Room_0, Office_2, and Office_3 from the Replica dataset [straub2019replica]. Employing the NeRF prior, our method effectively interpolates the unobserved geometry obscured by obstacles in the room, contrasting with Gaussian-based approaches that exhibit significant gaps, drastically compromising the completeness of the reconstruction.
TABLE V: The performance of ATE RMSE (cm) on the TUM RGB-D dataset [sturm2012evaluating]. For our method, we provide results both with and without Bundle Adjustment (BA).
@@
ORB Photo-SLAM [huang2024photo] 2.61 0.35 1.00 1.32
RTG-SLAM [peng2024rtg] 1.66 0.38 1.13 1.06
ESLAM [johari2023eslam] 2.30 1.10 2.40 2.00
Co-SLAM [wang2023co] 2.70 1.90 2.60 2.40
Frame-to-frame Loopy-SLAM [liso2024loopy] 3.79 1.62 3.41 2.94
SplaTAM [keetha2024splatam] 3.35 1.24 5.16 3.25
MonoGS [matsuki2024gaussian] 1.52 1.58 1.65 1.58
@@
Evaluations on TUM RGB-D Dataset We also conducted experiments on the TUM RGB-D dataset [sturm2012evaluating] and compared the ATE RMSE (cm) with both feature-based and frame-to-frame tracking systems in Table V. Our system achieves performance comparable to current state-of-the-art methods.